# On Creativity and Open-Endedness

### v2

@@ Abstract
ALife (ALife) as an interdisciplinary field draws inspiration and influence from a variety of perspectives. Scientific progress crucially depends, then, on concerted efforts to invite cross-disciplinary dialogue. The goal of this paper is to revitalize discussions of potential connections between the fields of CC (CC) and ALife, focusing specifically on the concept of OE (OE); the primary goal of CC is to endow artificial systems with creativity, and ALife has dedicated much research effort into studying and synthesizing OE and artificial innovation. However, despite the close proximity of these concepts, their use so far remains confined to their respective communities, and their relationship is largely unclear. We provide historical context for research in both domains, and review the limited work connecting research on creativity and OE explicitly. We then highlight specific questions to be investigated in future work, with the eventual goals of (i) decreasing conceptual ambiguity by highlighting similarities and differences between the concepts of OE, (ii) identifying synergy effects of a research agenda that encompasses both OE and creativity, and (iii) establishing a dialogue between ALife and CC research.
@@ Introduction
While our goal is ultimately to deepen our understanding of each concept and the requirements to reproduce them in artificial systems, we do not pretend to resolve these connections conclusively given limited space, time, and expertise. Instead, we ask provoking questions on the nature and constituents of these connections to establish a starting point for additional deep, dedicated interdisciplinary investigations.
By doing so, we continue and revitalize a prior discussion by Soros and Stanley [2016], asking “Are there nontrivial differences between creativity and OEE, or are they fundamentally the same concept?” Our treatment is more extensive and general, as it also considers OE beyond OEE (OEE). Existing work in CC focuses on OE as a property of tasks / exercises / problems [e.g. Epstein, 2011] or environments [e.g. Guckelsberger et al., 2016], but not as a characteristic of the process or the resulting product which ALife seeks to reproduce. Vice versa, ALife researchers have drawn connections from OE to creativity, but articulated only as sidelines. Instances of work spanning both fields and research agendas [Boden, 2009; Soros and Stanley, 2016] are promising, but few and isolated.
While this paper was written primarily with the goals and understanding of CC and ALife researchers in mind, creativity and OE are becoming dramatically more relevant beyond the boundaries of these communities, especially for AI more broadly [e.g. Hughes et al., 2024]. This paper is written for a wider audience, and we hope that our insights can inspire and advance research in AI and beyond.
@@
After nearly two decades of work on the problem, Packard et al. [2019] note that evolution’s dynamic production of diversity “has led life’s creative productivity to be called Open-Ended Evolution (OEE) in the field” [^2] The curious reader is directed towards Packard et al. [2019] for an excellent and more extensive introduction to OEE (OEE) and OE (OE).. The typical approach towards understanding and synthesizing the phenomenon has been to create virtual worlds imbued with features of the natural world thought to be key to unbounded generation (of whatever things, be it code snippets or embodied agents), and then performing systemic experiments probing exactly how long the generative process lasts and for what reasons it might fail. Notable systems establishing the foundation for current research include Tierra [Ray, 1992], Avida [Ofria and Wilke, 2004], PolyWorld [Yaeger, 1994], Geb [Channon and Damper, 2000], Cosmos [Taylor, 2004], Division Blocks [Spector et al., 2007], Evosphere [Miconi and Channon, 2005], and Chromaria [Soros and Stanley, 2014]. It should be noted that this list is not exhaustive.
Throughout these decades, many definitions of OEE (but, to the best of our knowledge, no definitions of OE) have been proposed. To only name a few, it has been defined as:
@@
Some of the ambiguity in conceptualisation and measurement may not be symmetrical between ALife and CC, offering an opportunity to mitigate these challenges in ALife, and similarly address open questions in CC. To exploit these opportunities though, we must better understand the connection between creativity and OE/OEE. This need is also demonstrated by the undifferentiated mentioning of creativity in writing about [e.g. Packard et al., 2019] and definitions of OE/OEE [e.g. Bedau, 1998], also quoted in this section.
@@
As our first contribution, we provide an up-to-date review of connections between OE and creativity previously drawn in related work. First, though, we highlight the scope of this review and which work has been excluded.
@@
Boden [2015a] has established an early conceptual connection between creativity and ALife. First, “psychological creativity” (involving human thought and activity) is contrasted with “biological creativity”, which is described as “the ability to generate new cells, organs, organisms, or species” and includes “the (phylogenetic) creativity of evolution, the (ontogenetic) creativity of morphogenesis, and the (autopoietic) creation of individual cells”. She then re-applies one of her most well known distinctions of creative mechanisms [Boden, 2004]: “the different sorts of surprise that we experience on encountering a creative idea correspond to different mechanisms for producing novelty. Those mechanisms mark three kinds of creativity: combinational, exploratory, and transformational” [Boden, 2015a]. More specifically:
@@
Closely related, Taylor [2019] relabels and connects different kinds of novelties defined by Banzhaf et al. [2016] with Boden’s types of creativity. Banzhaf et al. distinguish three types of novelty based on whether they require changes to a system’s model or meta-model. The model describes the system’s behavior through concepts which are themselves defined in a meta-model. Exploratory novelties (Banzhaf et al.: variation) result from applying the current dynamics of a system (summarised as a model) to multiple possible states. Expansive novelties (Banzhaf et al.: innovation) result from applying similar models that share concepts present in the current meta-model. Lastly, transformational novelties (Banzhaf et al.: emergence) can only result from models made of entirely new concepts, which necessitates a change in the meta-model description of the system’s dynamics. In a footnote, Taylor [2019] notes that these three types “fit closely with Boden’s concepts of exploratory, combinational and transformational creativity”. He does not elaborate on this further, except for saying that the term compositional is preferred over Boden’s combinational “to emphasise that the size of structures may increase, and that the specific arrangement and connections between components might be important”. Crucially, Taylor interprets the three types of novelties as flavors of ongoing adaptive novelty as a hallmark of OEE [Taylor et al., 2016, also mentioned above] and consequently considers them three types of open-ended processes. He thus draws parallels between Boden’s theory and ongoing adaptive novelty only. We note that this marks a distinction to Soros and Stanley’s [2016] mapping mentioned above which also includes ongoing adaptive complexity. This ambiguity has so far not been resolved.
@@
A long history of exchange between ALife and CC [Bown and McCormack, 2009; Saunders, 2012; Bown, 2012; McCormack, 2012; Guckelsberger et al., 2017; McCormack, 2019] bears witness of this endeavour’s inspirational and transformational potential. The very positive feedback and acceptance at one of the primary venues of ALife research highlight the willingness of the ALife community to take this initiative forward; separate calls by CC researchers to bridge communities [e.g. Cook and Colton, 2018] support that it will also fall on fertile CC ground.
@@ Acknowledgements
We thank our anonymous reviewers for their excellent feedback and support for this paper’s mission. LS, AA, and OW were supported by Cross Compass, Ltd.

### v4

@@ Abstract
ALife (ALife) as an interdisciplinary field draws inspiration and influence from a variety of perspectives. Scientific progress crucially depends, then, on concerted efforts to invite cross-disciplinary dialogue. The goal of this paper is to revitalize discussions of potential connections between the fields of CC (CC) and ALife, focusing specifically on the concept of OE (OE); the primary goal of CC is to endow artificial systems with creativity, and ALife has dedicated much research effort into studying and synthesizing OE and artificial innovation. However, despite the close proximity of these concepts, their use so far remains confined to their respective communities, and their relationship is largely unclear. We provide historical context for research in both domains, and review the limited work connecting research on creativity and OE explicitly. We then highlight specific questions to be investigated in future work, with the eventual goals of (i) decreasing conceptual ambiguity by highlighting similarities and differences between the concepts of OE and creativity, (ii) identifying synergy effects of a research agenda that encompasses both concepts, and (iii) establishing a dialogue between ALife and CC research.
@@ Introduction
While our goal is ultimately to deepen our understanding of each concept and the requirements to reproduce them in artificial systems, we do not pretend to resolve these connections conclusively given limited space, time, and expertise. Instead, we ask provoking questions on the nature and constituents of these connections as starting points for additional deep, dedicated interdisciplinary investigations.
By doing so, we revitalize a prior discussion by Soros and Stanley [2016], asking “Are there nontrivial differences between creativity and OEE, or are they fundamentally the same concept?” Our treatment is more extensive and general, as it also considers OE beyond OEE (OEE). Existing work in CC focuses on OE as a property of tasks / exercises / problems [e.g. Epstein, 2011] or environments [e.g. Guckelsberger et al., 2016], but not as a characteristic of the process or the resulting product which ALife seeks to reproduce. Vice versa, ALife researchers have drawn connections from OE to creativity, but articulated only as sidelines. Instances of work spanning both fields and research agendas [Boden, 2009; Soros and Stanley, 2016] are promising, but few and isolated.
In a more general sense, this paper continues a long-lasting tradition of knowledge exchange between the fields of ALife and CC. While it was written primarily with the goals and understanding of CC and ALife researchers in mind, creativity and OE are becoming dramatically more relevant beyond the boundaries of these communities, especially for AI more broadly [e.g. Hughes et al., 2024]. This paper is written for a wider audience, and we hope that our insights can inspire and advance research in AI and beyond.
@@
After nearly two decades of work on the problem, Packard et al. [2019] note that evolution’s dynamic production of diversity “has led life’s creative productivity to be called Open-Ended Evolution (OEE) in the field” [^2] The curious reader is directed towards Packard et al. [2019] for an excellent and more extensive introduction to OEE (OEE) and OE (OE).. The typical approach towards understanding and synthesizing the phenomenon has been to create virtual worlds imbued with features of the natural world thought to be key to unbounded generation (of whatever things, be it code snippets or embodied agents), and then performing systemic experiments probing exactly how long the generative process lasts and for what reasons it might fail. Notable systems establishing the foundation for current research include Tierra [Ray, 1996], Avida [Ofria and Wilke, 2004], PolyWorld [Yaeger, 1994], Geb [Channon and Damper, 2000], Cosmos [Taylor, 1999], Division Blocks [Spector et al., 2007], Evosphere [Miconi and Channon, 2005], and Chromaria [Soros and Stanley, 2014]. It should be noted that this list is not exhaustive.
Throughout these decades, many definitions of OEE have been proposed. To only name a few, it has been defined as:
@@
Some of the ambiguity in conceptualization and measurement may not be symmetrical between ALife and CC, offering an opportunity to mitigate these challenges in ALife, and similarly address open questions in CC. To exploit these opportunities though, we must better understand the connection between creativity and OE/OEE. This need is also demonstrated by the undifferentiated mentioning of creativity in writing about [e.g. Packard et al., 2019] and definitions of OE/OEE [e.g. Bedau, 1998], also quoted in this section.
@@
As our first contribution, we provide an up-to-date review of connections between OE and creativity previously drawn in related work, in chronological order. First, though, we clarify the review’s scope and which work has been excluded.
@@
To the best of our knowledge, Taylor was first to draw connections between creativity and OE already in 2002. Following the observation that “most [previous evolutionary ALife systems] are only capable of producing innovations of the ‘more-of-the-same’ variety’, he distinguishes such “basic” from “creative” OEE. Here, creativity is used as a qualifier of open-ended processes in evolution capable of “fundamental novelty” such as the “ability of individuals to interact with their (biotic and abiotic) environment with few restrictions, and to evolve mechanisms for sensing new aspects of this environment and for interacting with it in new ways”. Taylor highlights that creativity is distinct from OE, and that “a system capable of open-ended evolution is not necessarily creative”. He then elaborates on system design decisions shaping the potential for such fundamental novelty to emerge, comprising the design of the environment, interactions between individuals, and the individuals themselves. By doing so, he implicitly draws another connection to creativity, in that systems theories of creativity rely on the same elements [Taylor, 2002]. Crucially, Taylor uses creativity to emphasise fundamental novelty but does not explicitly discuss the value component of creativity which is considered essential in most definitions [Runco and Jaeger, 2012].
More than a decade later, Boden [2015a] established further conceptual connections between creativity and ALife more generally. We include this contribution as it influences the later discussion on connections to OE/OEE specifically. First, Boden distinguishes “psychological creativity” (involving human thought and activity) is contrasted with “biological creativity”, which is described as “the ability to generate new cells, organs, organisms, or species” and includes “the (phylogenetic) creativity of evolution, the (ontogenetic) creativity of morphogenesis, and the (autopoietic) creation of individual cells”. She then re-applies one of her most well known distinctions of creative mechanisms [Boden, 2004]: “the different sorts of surprise that we experience on encountering a creative idea correspond to different mechanisms for producing novelty. Those mechanisms mark three kinds of creativity: combinational, exploratory, and transformational” [Boden, 2015a]. More specifically:
@@
Closely related, Taylor [2019] relabels and connects different kinds of novelties defined by Banzhaf et al. [2016] with Boden’s types of creativity. Banzhaf et al. distinguish three types of novelty based on whether they require changes to a system’s model or meta-model. The model describes the system’s behavior through concepts which are themselves defined in a meta-model. Exploratory novelties (Banzhaf et al.: variation) result from applying the current dynamics of a system (summarised as a model) to multiple possible states. Expansive novelties (Banzhaf et al.: innovation) result from applying similar models that share concepts present in the current meta-model. Lastly, transformational novelties (Banzhaf et al.: emergence) can only result from models made of entirely new concepts, which necessitates a change in the meta-model description of the system’s dynamics. In a footnote, Taylor [2019] notes that these three types “fit closely with Boden’s concepts of exploratory, combinational and transformational creativity”. He does not elaborate on this further, except for saying that the term compositional is preferred over Boden’s combinational “to emphasise that the size of structures may increase, and that the specific arrangement and connections between components might be important”. Connecting to his pioneering 2002 work, expansive and transformational innovations map to what Taylor referred to as “creative”, and exploratory innovations to “basic” OEE [Taylor, 2024]. Crucially, Taylor [2019] interprets the three types of novelties exclusively as flavors of ongoing adaptive novelty as a hallmark of OEE [Taylor et al., 2016, also mentioned above]. This marks a distinction to Soros and Stanley’s [2016] earlier mapping which also includes ongoing adaptive complexity. This ambiguity has so far not been resolved.
@@
We note that ALife and CC researchers can look back at a long history of intellectual exchange. One of the earliest examples is Bentley and Corne’s 2002 collection “Creative Evolutionary Systems” [Bentley and Corne, 2002], followed by a landmark Dagstuhl meeting in 2009 [Boden et al., 2009] and a proliferation of cross-disciplinary work [e.g. Bown and McCormack, 2009; Bown, 2012; McCormack, 2012; Guckelsberger et al., 2017; McCormack, 2019]. This history bears witness of our endeavour’s inspirational and transformational potential. The very positive feedback and acceptance at this conference, one of the primary venues of ALife research, highlight the willingness of the ALife community to take this initiative forward; separate calls by CC researchers to bridge communities [e.g. Cook and Colton, 2018] support that it will also fall on fertile CC ground.
@@ Acknowledgements
We thank our anonymous reviewers for their excellent feedback and support for this paper’s mission, and Tim Taylor for further historical references included in this final version. LS, AA, and OW were supported by Cross Compass, Ltd.